OpenStudy (anonymous):

integrate ln(2x+1)dx

OpenStudy (anonymous):

We did this once?

OpenStudy (anonymous):

when?

OpenStudy (anonymous):

maybe i was not here

OpenStudy (anonymous):

Let p=2x+1 Then dp=2dx Rewrite your integral as: \[\frac{1}{2} \int\limits \ln(p)dp\] Then use integration by parts: u=ln(p) dv=dp du=1/p v=p so you have: \[\frac{1}{2}p \ln(p)-\frac{1}{2} \int\limits \frac{p}{p}dp=\frac{1}{2}p \ln(p)-\frac{1}{2}p+C\] Then replacing p you have: \[\frac{1}{2}(2x+1)\ln(2x+1)-\frac{1}{2}(2x+1)+C\]